OpenStudy (anonymous):

Which statements are correct interpretations of this graph? Choose exactly two answers that are correct. A. 3 pages are edited every 5 min. B. 5 pages are edited every 3 min. C. of a page is edited per minute D. pages are edited per minute

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

The answer is A. Look at the graph 5 minutes pass, 3 pages are edited 10 minutes pass, 6 pages are edited 15 minutes pass, 9 pages are edited Minutes go up by 5, Pages go up by 3.

OpenStudy (anonymous):

Its MC there are two answers

OpenStudy (anonymous):

c = 6/10 and d = 1 2/3

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Oh, well 0.6 pages are edited every minute because: 3 / 5 = 0.6

OpenStudy (anonymous):

thx